### v1

# Examining Differential Item Functioning (DIF) in Self-Reported Health Survey Data: Via Multilevel Modeling

Abstract
Few health-related constructs or measures have received critical evaluation in terms of measurement equivalence, such as self-reported health survey data. Differential item functioning (DIF) analysis is crucial for evaluating measurement equivalence in self-reported health surveys, which are often hierarchical in structure. While traditional DIF methods rely on single-level models, multilevel models offer a more suitable alternative for analyzing such data. In this article, we highlight the advantages of multilevel modeling in DIF analysis and demonstrate how to apply the DIF framework to self-reported health survey data using multilevel models. For demonstration, we analyze DIF associated with population density on the probability to answer “Yes” to a survey question on depression and reveal that multilevel models achieve better fit and account for more variance compared to single-level models. This article is expected to increase awareness of the usefulness of multilevel modeling for DIF analysis and assist healthcare researchers and practitioners in improving the understanding of self-reported health survey data validity.
keywords:
differential item functioning, measurement equivalence, multilevel modeling, health disparity, population density, depression
1 Introduction
To examine the disproportionate impact of health interventions or other determinants on health outcomes for certain populations, including disadvantaged minorities, it is crucial to establish measurement equivalence when evaluating outcomes for different demographic groups. However, few health constructs or measures have undergone critical evaluation in this regard [1], despite the routine practice of validating psychometric instruments in research. Recently, more health studies have employed differential item functioning (DIF) analysis to examine measurement equivalence and the differential impact of groups differing in gender and ethnicity [e.g., 2, 3, 4, 5, 6]. DIF analysis assesses whether there is a systematic difference in the probability of obtaining a targeted response between groups while controlling for a latent trait [7]. That said, some health studies did not clearly describe the method used for DIF analysis [e.g., 3], and conventional DIF methods based on single-level models were frequently used to analyze health survey data with a nesting structure [e.g., 2], for which recent DIF methods based on multilevel models are better alternatives in terms of accuracy of DIF estimation.
2 Challenge: Examine DIF in self-reported health survey data
Self-reported health survey data encapsulate rich information about physical, mental, and social health trends of the studied population [e.g., 8, 9]. However, its inherent subjectivity makes it susceptible to geographical, cultural, and socioeconomic biases, which can threaten the validity of the data and pose challenges in getting to know the truth [e.g., 10]. The utilization of self-reported health survey data in public health research and interventions necessitates a rigorous examination of the validity of this data. It underscores the importance of adopting appropriate methods that can navigate the complexities associated with self-reported health data and improve the accuracy of our findings.
Existing research has shown that bias can manifest itself as construct bias, method bias, and item bias [e.g., 11, 12]. The framework of differential item functioning [DIF; 13] underpins many modern methods to investigate item bias, specifically. In the psychometric literature, DIF analysis examines whether there is a systematic between-group difference in examinees’ probability to answer an item correctly while controlling for their ability [7]. While DIF is not equivalent to item bias, the former informs judgment on the latter when its presence is attributed to unintended content in the studied item [14, 15].
To examine systematic between-group differences in self-reported health survey data, we can apply existing DIF methods by generalizing the DIF concept with the following modifications:
  1. 1.
    focusing on the response of interest instead of a “correct” response,
  2. 2.
    describing the subjects who provide responses as “respondents” instead of “examinees,”
  3. 3.
    using a latent trait variable instead of “ability,” selecting one closely related to the probability of obtaining the response of interest (similar to how “ability” relates to the probability of a correct response in the conventional DIF framework).
Some existing DIF methods examine both uniform and nonuniform DIF while many others do not. Uniform DIF occurs when respondents from one group consistently outperform respondents of the same latent trait from another group; nonuniform DIF exists when this performance difference is not consistent [16].
Most DIF methods currently used in health research, as well as in the broader DIF literature, are based on single-level modeling [e.g., 2]. They are not best suited for self-reported health survey data that has a nesting or hierarchical structure (e.g., individual responses are nested within larger units differentiated by population density, economy, and culture). These single-level DIF methods include the Mantel-Haenszel procedure [MH; 17], the logistic regression procedure [LR; 18], Lord’s Wald test [19], and the IRT likelihood-ratio test [IRT-LR; 20]. The first two do not require the IRT model and are labeled with “the observed-score approach,” whereas the latter two require the IRT model and are categorized as “the IRT approach” [21, 22]. An extensive review of these methods can be found in Chen [23].
Single-level modeling approaches have at least three problems when the response data has a hierarchical structure. First, they waste information and render inaccurate estimates because of the lack of consideration of variances between hierarchies of the data [24]. Second, they do not help model the effect of higher-level variables associated with DIF on the outcome variable at the lower level [25]. Third, their model estimates are limited to describing the sample and are not generalizable to describe the population [26]. Moreover, the single-level modeling for DIF analysis is limited to the analysis of two groups at a time, not suitable for DIF analyses involving more than two groups [23].
3 Solution: Multilevel modeling for DIF analysis
For the self-reported health survey data that is hierarchical in nature, multilevel modeling has been demonstrated to be a superior alternative to single-level modeling. It effectively addresses issues that arise in differential item functioning (DIF) methods built on single-level models. First, multilevel modeling is advantageous when fundamental assumptions in single-level modeling such as independence and homoscedasticity are violated, as is often the case with hierarchical data [24]. In this regard, multilevel modeling can account for correlated errors resulting from the dependency of individual observations within a nested structure, and it does not require the same sample size or variance for higher-level units due to its randomization approach [27].
Second, multilevel modeling allows for the assessment of higher-level variables’ impact on DIF [25]. For instance, it can evaluate how cluster-level factors, like access to healthcare facilities, influence individual responses in health surveys, and identify whether these cluster characteristics contribute to DIF observed between population groups at the lower level of the model. Third, results from multilevel modeling are generalizable to the population because it incorporates random effects from higher-level clusters [26]. This ensures that conclusions about DIF in self-reported response data are applicable to a broader range of regions. Fourth, multilevel modeling accommodates multi-group analyses by using multiple dummy-coded variables to cover more than two groups [28]. For instance, it can include multiple age groups in health data, allowing researchers to analyze and compare health outcomes across various age categories while accounting for cluster-specific random effects.
The first published work on DIF using multilevel modeling was by Swanson et al. [28], who adopted a multilevel logistic regression model. This approach significantly improved the accuracy of DIF estimates and pooled information across items to explain sources of DIF, when including random item effects. Subsequently, den Noortgate and de Boeck [29] proposed a series of multilevel models for DIF analysis, grounded in the Rasch model within item response theory (IRT). These models demonstrated the potential to integrate random effects associated with respondents, items, or other factors, along with item- or group-related covariates, to explain DIF. French and Finch [30] proposed a multilevel MH statistic that adjusts the traditional MH statistic using a ratio of random variances from the multilevel model from Swanson et al. [28]. Huang and Valdivia [31] introduced a multilevel Wald test for polytomous items based on Lord’s Wald $\chi^{2}$ test and a two-stage procedure about first identifying anchor items and then evaluating DIF items for non-anchor items.
In the following sections, we focus on the multilevel logistic regression procedure from Swanson et al. [28] because of its versatility. Compared to the models from den Noortgate and de Boeck [29] and from Huang and Valdivia [31], the multilevel logistic regression procedure does not require IRT or its stringent assumptions, such as local independence of item responses and monotonicity of the probability of the targeted response [32]. Meanwhile, to enhance the utility of the multilevel model in Swanson et al. [28], we have adapted it into a more general version that permits the inclusion of various types of random effects. For clarity, we also cover its single-level counterpart, the logistic regression procedure [18], to help readers understand how the multilevel logistic regression evolved from its origins in single-level modeling. Hereafter, we refer to the multilevel logistic regression procedure as “multilevel LR” and the logistic regression procedure as “traditional LR.” Also, we use the word “group” when referring to the group membership variable for group comparisons in DIF analysis, and the word “cluster” when referring to units of analyses across levels in the hierarchical data structure in multilevel modeling.
3.1 Multilevel LR
The multilevel LR [28] was initially designed to include random effects associated with items, accounting for variations among items and offering more accurate DIF estimates compared to the traditional LR [18]. We have modified its notation to allow the incorporation of random effects from higher-level clusters in hierarchical data (e.g., access to healthcare, geographical locations, age groups, education levels). In this modified version, the level-1 submodel pertains to responses to a studied item from individual respondents (indexed by $p$) within a level-2 stratum (indexed by $i$), similar to the binary logistic regression model used in traditional LR but without an interaction term. This level-1 submodel is expressed as
$\mbox{logit}(\pi_{pi})=\beta_{0i}+\beta_{1i}\theta_{p}+\beta_{2i}g_{p},$ (1)
where
$\pi_{pi}$ is the probability for respondent $p$ to provide the targeted response to the studied item (i.e., the “probability targeted”);
$\beta_{0i}$ denotes the log-odds of the “probability targeted” in the reference group;
$\beta_{1i}$ is the effect of respondents’ latent trait on the log-odds of the “probability targeted” in the reference group;
$\beta_{2i}$ is the effect of respondents being in the reference group on the log-odds of the “probability targeted;”
$\theta_{p}$ is the latent trait of respondent $p$ (e.g., it could be proxied by a continuous variable concerning the respondent’s overall health status);
$g_{p}$ is a dummy variable representing the group membership of respondent $p$ ($g_{p}=0$ for the reference group, $g_{p}=1$ for each focal group).
The level-2 submodel in the multilevel LR accounts for random effects associated with higher-level clusters, written as
$\begin{array}[]{lll}\beta_{0i}&=&\delta_{00}+U_{0i},\\ \beta_{1i}&=&\delta_{10}+U_{1i},\\ \beta_{2i}&=&\delta_{20}+\delta_{21}v_{1}+\delta_{22}v_{2}+\cdots+\delta_{2V}v_{V}+U_{2i},\end{array}$
where
$\delta_{00}$ is the average log-odds of the “probability targeted;”
$\delta_{10}$ is the average effect associated with respondents’ latent trait;
$\delta_{20}$ is the average effect associated with being in the focal group;
$\delta_{21}$, $\delta_{22}$, $\cdots$, $\delta_{2V}$ denote the effect of covariates $v_{1}$, $v_{2}$, $\cdots$, $v_{V}$ (either continuous or dummy-coded) that are associated with level-2 clusters; they are useful to explain why DIF occurs, but can be omitted if no level-2 covariate is investigated here;
$U_{0i}$ is the random effect associated with level-2 clusters on the intercept $\beta_{0i}$;
$U_{1i}$ and $U_{2i}$ denote the random effect associated with level-2 clusters on the slopes $\beta_{1i}$ and $\beta_{2i}$, respectively;
$U_{0i}$, $U_{1i}$, and $U_{2i}$ follow a multivariate normal distribution with the mean being 0.
Using this multilevel LR, we can conclude that the studied item exhibits DIF if $\beta_{2i}$ is significantly different from 0.
3.2 Traditional LR
In the traditioinal LR proposed by Swaminathan and Rogers [18], the binary logistic regression model used for DIF analysis specifies the effect of respondents’ latent trait and group membership on the log-odds of the probability of the targeted response to a studied item. In Swaminathan and Rogers [18], the authors first rewrote the DIF model inherent in the widely used Mantel-Haenszel procedure using binary logistic regression:
$\mbox{logit}(\pi)=\beta_{0}+\sum_{k=1}^{K}\beta_{k}\theta_{k}+\tau g,$ (2)
where
$\pi$ is the probability to provide the targeted response to the studied item;
$\beta_{0}$ denotes the average log-odds of the “probability targeted” in the reference group;
$\theta_{k}$ is the $k$th total score (in discrete units) in a fixed-item instrument for respondents;
$\tau$ is equivalent to the common odds ratio in the Mantel-Haenszel procedure;
$g$ is the group membership ($g$ = 0 for the reference group, $g$ = 1 for the focal group) of respondents.
Inspired by Mellenbergh [16], Swaminathan and Rogers [18] revised (2) by specifying separate terms for uniform and nonuniform DIF, written as
$\mbox{logit}(\pi)=\beta_{0}+\beta_{1}\theta+\beta_{2}g+\beta_{3}(\theta g),$ (3)
where
$\theta$ denotes respondents’ latent trait (proxied by the examinee’s total score);
$\beta_{1}$ is the effect of respondents’ latent trait on the log-odds of the “probability targeted” in the reference group;
$\beta_{2}$ is the effect of respondents being in the reference group on the log-odds of the “probability targeted;”
$\beta_{3}$ is the ability-moderated group difference in the effect on the log-odds of the “probability targeted;”
$g$ is respondents’ group membership ($g$ = 0 for the reference group, $g$ = 1 for the focal group).
Using the traditioinal LR specified in (3), we can conclude that the studied item exhibits nonuniform DIF if $\beta_{3}$ is significantly different from 0, and it exhibits uniform DIF if $\beta_{3}$ is not significantly different from 0 whereas $\beta_{2}$ is.
The traditional LR was established for analyzing two groups at a time. While a common practice in psychometrics is to perform pairwise comparisons using the traditional LR [e.g., 33], this approach can be computationally intensive and lead to increased Type-I error rates and reduced power [34, 35]. In contrast, the multilevel LR allows for multi-group analyses by employing multiple dummy variables to represent more than two groups [28], while controlling for cluster-specific random effects to ensure accuracy and generalizability.
4 Example: Analysis of population-density DIF in NSDUH data
To show the impact of multilevel modeling on DIF results, below we describe our DIF analyses when using both the multilevel LR and the traditional LR for comparison. The NSDUH data released in 2022 was utilized in these analyses. It provides comprehensive information on substance use and mental health trends among the U.S. population aged 12 and older. We are interested in examining DIF associated with population density because existing research has demonstrated that population density can affect the quality of life, health perceptions, and reported behaviors [36, 37, 38, 39]. In our multilevel analysis, we utilize a two-level model in which the level-2 clusters are defined by respondents’ education levels, as existing research has unveiled that an additional year of education could reduce the likelihood of reporting depression and anxiety [40].
Specifically, we want to examine whether DIF exists for respondents from high and low population density areas in the probability to answer “Yes” to a selected NSDUH item. This item asks respondents the following question: “Have you ever in your life had a period of time lasting several days or longer when most of the day you felt sad, empty or depressed?” The latent trait variable to be controlled for is a score assigned to respondents that reflects respondents’ psychological distress level over the past 30 days, based on their responses to another six questions. The reference group is the respondents from areas with low population density, and the focal group is the respondents from areas with high population density. The variable describing respondents’ education level is utilized for accounting for variations between the higher-level clusters. We acknowledge the potential for DIF in this latent trait variable originally derived from six survey items, but we do not examine DIF for this variable to focus on demonstrating the application of the multilevel LR rather than conducting a comprehensive DIF analysis across all measures.
4.1 Data preparation
The original NSDUH data file contains 2,605 variables and 59,069 respondents, composed by 27,047 males and 32,022 females aged 12 and older. We keep the following four variables for DIF analysis:
  1. 1.
    The item response variable (named “ADDPREV” in the survey). It contains the following responses to the selected survey item. It is recoded to have 1 for the response “Yes” and 0 for “No.”
  2. 2.
    The group membership variable (named “PDEN10”). It indicates the population density of the area where respondents are residing. It is recoded to have 1 for “high-density” (i.e., with one million or more people) and 0 for “low-density” (i.e., with fewer than one million people), based on Core Based Statistical Area (CBSA) classifications. These areas are also known as “large metropolitan” and “small metropolitan” areas [41].
  3. 3.
    The latent trait variable (named “KSSLR6MON”). It is a score that describes respondents’ distress, ranging from 0 to 24. This score is closely associated with the probability of obtaining the response of interest in the item response variable.
  4. 4.
    The level-2 variable (named “IREDUHIGHST2”). It describes respondents’ education level, and there are 11 distinct levels, ranging from 1 for “completing fifth grade or less” to 11 for “completing college or more advanced education.”
Appendix C displays information from the NSDUH codebook about how the selected categorical variables were originally coded.
Table 1: Sample size for each data cluster.
Level-2 clusters Group Response
Edu. level Low-density High-density No Yes
1 126 140 224 42
2 75 115 172 18
3 45 49 76 18
4 158 102 203 57
5 314 206 391 129
6 407 273 488 192
7 1228 1059 1590 697
8 6488 4430 7435 3483
9 5208 3895 5323 3780
10 2286 1563 2463 1386
11 7630 7874 9780 5724
Considering that the multilevel LR and the traditional LR use the binary logistic regression that are limited to the binary response data analysis, we focus on the respondents associated with “Yes” and “No,” removing the respondents that answer “Don’t know,” refuse to answer or leave a blank, or are led to skip the selected survey item due to the survey logic. So, for further analysis, we first drop 13,242 respondents associated with neither “Yes” nor “No.” Then, because we focus on DIF associated with population density, we exclude 2,156 respondents associated with non-CBSA classifications for which there is a lack of information of the population density. In this way, we retain 43,671 eligible respondents. They comprise 19,348 males and 24,323 females, with no missing values. Table 1 displays sample sizes for the level-2 clusters, which are large enough for multilevel modeling [42, 43].
4.2 Analysis and Results
We carried out DIF analyses using a series of models: three multilevel models associated with the multilevel LR shown in (1), and the single-level traditional LR model shown in (3). The first multilevel model we utilize does not contain explanatory variables in the level-1 submodel, which makes it the commonly known null model or empty model that sets a baseline for understanding the proportion of variance explained by the hierarchical structure of the data [27]. The second multilevel model is the exact model used in the multilevel LR, incorporating both the latent trait and group membership variables in the level-1 submodel. The third multilevel model adds the interaction term between the latent trait and group membership, the same as the interaction term in (3), to check if the model performance would be improved when adding this interaction. These three multilevel models are named “Model 0,” “Model 1,” and “Model 2.” The single-level model is labeled with “Model 3.”
The modeling results in Table 2 reveal several key insights regarding the performance of different models on DIF associated with population density in responses to the selected NSDUH item about depression. The null model (Model 0) serves as a baseline, showing the proportion of variance attributable to the hierarchical structure of the data. Its adjusted intraclass correlation coefficient [ICC; 24] of 0.071 indicates that 7.1% of the response variability stemmed from differences between education levels. This highlights the importance of considering this structure in DIF analyses with this data. Model 1, incorporating both the latent trait and group membership variables, displays a significant negative effect of high population density on the probability of responding “Yes” to the item. This suggests that this item has DIF, as individuals residing in high-density areas were less likely to provide the targeted response to the selected item compared to those in low-density areas, after accounting for all these individuals’ latent trait about depression. Model 2, which includes the interaction term, does not show a better model fit compared to Model 1 in this specific example. This implies that the interaction term does not substantially affect the model’s performance. Model 3, the single-level model, does not reveal a significant effect of high population density, which leads to the conclusion that there is no DIF, contrasting with the result from Model 1. This difference underscores the importance of model selection when conducting DIF analyses.
Overall, Models 1-2 outperform others, based on fit statistics. Both Model 1 and Model 2 show a low intercept variance ($\tau_{0}^{2}$ = 0.22), indicating an advantage in capturing more data variability. Additionally, they share the lowest AIC, BIC, and -2LL values, and the highest $R^{2}$ values. Their $R^{2}$ values signify that each of the two models explains 37.4% of the data variability through random and fixed effects and 33.1% solely through fixed effects. In contrast, the single-level model, Model 3, explains 26.9% of the variance, which is less favorable.
Table 2: Results from a series of models.
Model 0 Model 1 Model 2 Model 3
Fixed effects
$\beta_{0i}$1.03∗∗∗ -2.39∗∗∗ -2.41∗∗ -1.90∗∗∗
(0.15) (0.15) (0.15) (0.03)
$\beta_{1i}$ 0.25∗∗∗ 0.25∗∗∗ 0.25∗∗∗
(0.00) (0.00) (0.00)
$\beta_{2i}$0.12∗∗∗ -0.09 -0.07
(0.02) (0.04) (0.04)
$\beta_{3i}$0.01 -0.00
(0.01) (0.01)
Variances
$\tau_{0}^{2}$ 0.25 0.22 0.22
$\sigma^{2}$ 3.29 3.29 3.29
N 43671 43671 43671 43671
AIC 56460 44197 44198 44697
BIC 56477 44232 44241 44732
-2LL 56455 44189 44188 44689
ICC 0.071 0.064 0.064
$R^{2}_{marg}$ 0.000 0.331 0.331 0.269
$R^{2}_{cond}$ 0.071 0.374 0.374
  • ${}^{*}p<0.05$; ${}^{**}p<0.01$; ${}^{***}p<0.001$.
Researchers might consider Model 2 when separate interpretations of uniform and nonuniform DIF are desired, analogous to the traditional LR approach. A significant interaction term in Model 2 suggests the presence of non-uniform DIF, where the effect of population density on the probability of responding “Yes” varies across the continuum of psychological distress scores. In contrast, a non-significant interaction term but a significant coefficient for the group membership variable would indicate uniform DIF, suggesting a constant difference in the probability of responding “Yes” between the high- and low-density groups regardless of the level of psychological distress.
It is important to acknowledge potential limitations of this analysis. While the sample size for each level-2 cluster appears sufficient based on existing recommendations, further exploration of the data distribution within each cluster could be beneficial [27]. This might involve examining potential skewness or outliers that could influence model estimates. Additionally, although we controlled for a latent trait variable reflecting psychological distress, the inclusion of other relevant covariates (e.g., socioeconomic status, access to healthcare) might provide a more comprehensive picture of potential DIF effects.
In conclusion, our results highlight the strengths of multilevel modeling for DIF analysis in health research with clustered data. By accounting for the hierarchical structure of the data (respondents nested within education levels), multilevel models provided a more accurate picture of the underlying relationships. This is evident in the superior model fit statistics (lower AIC, BIC, and -2LL values, and higher R-squared values) of multilevel models and the identification of DIF that was missed by the single-level model. Future research could explore the generalizability of our findings by applying multilevel DIF analysis to other self-reported health survey data.
5 Discussion
Multilevel modeling we describe in this article is a useful tool for examining DIF in the data from health-related constructs or measures, such as self-reported health survey data. By leveraging this approach, researchers can gain a more accurate understanding of DIF and glean valuable insights into health survey data, ultimately enhancing the precision and reliability of health research findings. Built on prior research that highlights the benefits of multilevel modeling in DIF analysis, our study provides a practical application of how this approach can be used with self-reported health survey data.
Given that single-level LR is not suitable for polytomous items or multiple groups and that we want to have a comparison between single-level and multilevel LR models, our empirical analysis centers on a dichotomous item and a variable involving two groups for DIF assessment. Rather than making a blanket recommendation in favor of multilevel models, we intend to raise awareness of how to use multilevel LR to analyze DIF and how to choose the appropriate model when a series of competing models are available. Different from some other studies that evaluated models via Monte Carlo simulation [e.g., 44], our empirical analysis examines a series of competing models based on model fit indices. Our findings indicate that multilevel modeling more effectively captures variability in our data compared to single-level modeling, which is crucial to accurate DIF detection.
Future research is crucial to comprehensively explore the cost-benefit balance of multilevel modeling for DIF analysis in dynamic health survey settings. Monte Carlo simulations can be a powerful tool to systematically evaluate model performance under a wide range of data structures. These simulations can vary factors such as ICC, number and size of clusters, and the magnitude and proportion of DIF items. Studies by Moineddin et al. [45] and [46] provide valuable groundwork, but further exploration is necessary. While initial efforts have been made to explore multilevel DIF analysis [e.g., 47, 48, 49], limitations exist regarding model construction, ICC calculation, and generalizability [23]. Scrutinizing model performance under various realistic scenarios that mimic real-world health survey data is essential. This includes exploring the impact of factors like missing data patterns (missing completely at random, missing at random, or missing not at random) and different types of DIF (uniform vs. non-uniform) on model performance. Also, researchers might consider developing or applying more advanced models capable of multilevel modeling for DIF analysis with polytomous items or more than two groups, ideally built on existing models without IRT-related assumptions such as ordinal logistic regression [50, 51] or nonparametric measures [e.g., 52]. By addressing these limitations, we can solidify the role of multilevel modeling in DIF analysis for health research. This will ultimately lead to more robust and reliable estimates captured through self-reported surveys.
Declarations
[…]
Funding The authors declare that no funds, grants, or other support were received during the preparation of this manuscript.
[…]
Competing Interests The authors have no relevant financial or non-financial interests to disclose.
[…]
Availability of Data and Materials The NSDUH data and the codebook used in this article can be downloaded here: https://www.datafiles.samhsa.gov/dataset/national-survey-drug-use-and-health-2022-nsduh-2022-ds0001
Appendix A R code
A.1 For data preparation
1# initialize
2load("NSDUH_2022.Rdata")
3d <- NSDUH_2022
4
5# keep relevant data
6d <- d[, c("addprev", # item
7 "PDEN10", # group membership
8 "KSSLR6MON", # latent trait
9 "IREDUHIGHST2" # Level-2 variable
10 )]
11
12# rename variables
13names(d) <- c("item", "group", "score", "level2")
14
15# remove invalid values
16d <- d[which(d\$group != 3), ] # non-CBSA areas
17d <- d[which(d\$item %inc(85, 94, 97, 98, 99) == FALSE),] # invalid response values
18
19# recode two variables
20d\$item[d\$item == 2] <- 0 # to have 1=Yes, 0=No
21d\$group[d\$group == 2] <- 0 # 1=high, 0=low
22
23# check missing values
24apply(d, MARGIN = 2, function(x){
25 table(x, useNA = "always")
26})
27
28# check level-2 sample sizes
29table(d\$level2, d\$group)
30table(d\$level2, d\$item)
31
32# finalize the data
33dta <- d
34dta\$group <- as.factor(dta\$group) # to make sure it is formated as a categorical variable, allowing glmer() to proceed with dummy coding
A.2 For DIF analyses
1library(lme4)
2
3# Model 0
4model0 <- glmer(item ~ 1 + (1 | level2), data = dta, family = binomial)
5summary(model0)
6tau00 <- as.data.frame(VarCorr(model0))[4]
7icc <- performance::icc(model0)
8resid <- tau00/icc\$ICC_adjusted - tau00
9#// residual variance
10
11# Model 1
12model1 <- glmer(item ~ score + group + (1 | level2), data = dta, family = binomial)
13summary(model1)
14tau00 <- as.data.frame(VarCorr(model1))[4]
15icc <- performance::icc(model1)
16resid <- tau00/icc\$ICC_adjusted - tau00
17
18# Model 2
19model2 <- glmer(item ~ score + group + score*group + (1 | level2), data = dta, family = binomial)
20summary(model2)
21tau00 <- as.data.frame(VarCorr(model2))[4]
22icc <- performance::icc(model2)
23resid <- tau00/icc\$ICC_adjusted - tau00
24
25# Model 3
26model3 <- glm(item ~ score + group + score*group, data = dta, family = "binomial")
27summary(model3)
28
29# compile results
30library(texreg)
31screenreg( list(model0, model1, model2, model3) )
32
33# obtain other statistics with this one
34sjPlot::tab_model(list (model0, model1, model2, model3))
Appendix B DIF Methods on Single-level Models
An extensive review of DIF methods can be found in Chen [23]. Table 3 summarizes the limitations of some popular DIF methods based on single-level models. In this appendix, we detail the Mantel-Haenszel procedure [MH; 17] and the Lord’s Wald test with worked examples, including the R code we used for computations. Our analyses were done via difR package [53] with R4.4.0 in RStudio.
Table 3: Limitations of the single-level DIF methods reviewed in Chen [23].
Method Limitations
MH [17] Limited to two groups; using observed, discrete total scores; assuming that scores are equally spaced, and that the common odds ratio remains constant.
LR [18] Limited to two groups; lacking a valid effect size measure; inflated errors with uniform DIF detection, with 3PL items.
Lord’s Wald test [19] Limited to two groups; item parameter estimates depend on the item set; lacking a valid effect size measure.
IRT-LR [20] Limited to two groups; computationally intensive; lacking a valid effect size measure.
B.1 The Mantel-Haenszel Procedure
The Mantel-Haenszel (MH) procedure, introduced by Holland and Thayer [17], is an extension of Pearson’s chi-square test of independence that uses odds ratios from tables to quantify the association between two dichotomous variables, conditional on a third variable that could be a confounder. In Holland’s work, a 2$\times$2 table is created to cross-classify the type of responses (i.e., correct vs. incorrect) by group (i.e., reference group vs. focal group) for the $k$th total score ($k$ = 1, 2, $\cdots$, $K$), demonstrated in Table 4. The row margins are fixed to 1 by design because an item response is either the targeted response (e.g., “Yes”) or not (e.g., “No”), and hence the cell values follow a binominal distribution for each row. The probabilities of answering an item correctly in the reference and focal groups are $\pi_{Rk}$ and $\pi_{Fk}$, respectively, conditional on the $k$th total score. Correspondingly, the probabilities of obtaining a targeted response in the reference and focal groups are $(1-\pi_{Rk})$ and $(1-\pi_{Fk})$, respectively.
Table 4: A 2$\times$2 table in probabilities (by group membership and item response) for the $k$th total score.
Yes (1) No (0) Total
Reference group $\pi_{Rk}$ $1-\pi_{Rk}$ 1
Focal group $\pi_{Fk}$ $1-\pi_{Fk}$ 1
In the MH, the null hypothesis is that there is no DIF, or in other words, that there is conditional independence between an examinee’s group membership and their response after accounting for their total scores. It is expressed as
$H_{0}\text{: }\dfrac{\pi_{Rk}}{1-\pi_{Rk}}=\dfrac{\pi_{Fk}}{1-\pi_{Fk}}\text{ for all }k.$
The alternative hypothesis is
$H_{1}\text{: }\dfrac{\pi_{Rk}}{1-\pi_{Rk}}=\alpha\dfrac{\pi_{Fk}}{1-\pi_{Fk}}\text{ for all }k,$
where $\alpha$ is the common odds ratio shared by the $k$ tables.
Probabilities in Table 4 can be expressed as counts, as shown in Table 5.
Table 5: A $2\times 2$ table in counts (by group and response) for the $k$th total score.
Yes (1) No (0)
Reference group $A_{k}$ $B_{k}$
Focal group $C_{k}$ $D_{k}$
Then the MH test statistic (with continuity correction to improve the approximations of significance levels) can be expressed as
$\chi^{2}_{MH}=\dfrac{(|\sum_{k}A_{k}-\sum_{k}E(A_{k})|-\dfrac{1}{2})^{2}}{\sum_{k}\mbox{var}(A_{k})},$
which follows a chi-square distribution, which assumes independence of observations.
Holland and Thayer [17] proposed an effect size measure of DIF, specified as
$\hat{\Delta}_{MH}=-2.35\ln(\hat{\alpha}),$
where the logarithm puts $\hat{\alpha}$ on an additively symmetric scale, making it convenient for interpretation. Since the 1990s, ETS has been flagging DIF items with $A$ (i.e., negligible/nonsignificant DIF), $B$ (i.e., slight to moderate DIF), and $C$ (i.e., moderate to large DIF) using $\chi^{2}_{MH}$ and $\hat{\Delta}_{MH}$, as documented in Zwick [54].
The MH method is simple and widely used in educational assessment development, but it has some limitations. Firstly, MH uses observed, discrete total scores as a proxy for a latent trait, resulting in information loss. Secondly, using discrete total scores as the third dimension for analyzing two-way tables involves unrealistic assumptions: that total scores are equally spaced and that the common odds ratio $\alpha$ remains constant for different total scores [54].
Application in R. Using the data of the respondents answering either “Yes” or “No” to the selected item we described in the main text, we found the MH test statistic for the item is 15.406, its $p$-value is 0.0001, which is highly significant, suggesting strong evidence against the null hypothesis that there is no DIF. The common odds ratio is greater than 1 ($\alpha_{MH}=1.0980$), suggesting that the respondents in the reference group is more likely to answer “Yes” to the selected item than the focal group.
However, it is important to consider both the statistical significance ($p$-value) and the effect size when interpreting these results, as a statistically significant finding may not always have practical significance. The effect size measure of DIF indicates a negligible effect ($\hat{\Delta}_{MH}=-0.2197$, labeled with $A$).
Table 6: Results from the Mantel-Haenszel procedure.
Measure Statistic $p$-value
$\chi^{2}_{MH}$ 15.406 ∗∗∗ 0.0001
$\alpha_{MH}$ 1.098
$\hat{\Delta}_{MH}$0.220
Effect Size $A$
  • ${}^{*}p<0.05$; ${}^{**}p<0.01$; ${}^{***}p<0.001$.
In summary, the MH test results indicate that DIF exists for the item. Nevertheless, its effect size is negligible, suggesting that the practical significance of this DIF might be small. The following R code demonstrates how to use an existing R function difMH() to do the MH analysis, as well as how to directly compute the MH test statistic without that function. We found that the MH test statistic is 15.406 via either of the procedures.
1# label the groups with "F" for reference groups and "R" for focal groups
2dta\$group <- ifelse(dta\$group == 1, "F", "R")
3
4# MH test (via an existing R function)
5library(difR)
6mh_result <- difMH(Data = as.data.frame(dta\$item), group = dta\$group, focal.name = ’F’, match = dta\$score)
7mh_result
8#// display results
9
10# MH test (via direct computation)
11l <- split(dta, as.factor(dta\$score))
12
13vals <- list()
14for(i in 1:length(l)){
15
16 x <- l[[i]]
17
18 d <- data.frame(table(x\$group, x\$item))
19
20 a <- d\$Freq[which(d\$Var1 == "R" & d\$Var2 == 1)]
21 n <- sum(d\$Freq)
22
23 m1 <- sum(d\$Freq[which(d\$Var1 == "R")])
24 n1 <- sum(d\$Freq[which(d\$Var2 == 1)])
25 a_expected <- m1*n1/n
26
27 m2 <- sum(d\$Freq[which(d\$Var1 == "F")])
28 n2 <- sum(d\$Freq[which(d\$Var2 == 0)])
29 variance <- (m1/10*m2/10*n1/10*n2/10)/(n^2*(n-1))*10^4
30 #// function: m1*m2*n1*n2/(n^2*(n-1)) -- we added "/10" and "*10^4" here to avoid the overflown error in R
31
32 t <- data.frame(cbind(a, a_expected, variance))
33 names(t) <- c("a", "a_expected", "variance")
34 vals[[i]] <- t
35}
36vals <- do.call(rbind, vals)
37
38chi <- (abs(sum(vals\$a) - sum(vals\$a_expected)) - 1/2)^2 / sum(vals\$variance)
39chi
40#// display the test statistic
B.2 Lord’s Wald Test
Lord’s Wald test [19], derived from the traditional Wald test [55], is historically significant. It requires the item response theory models (IRT) to estimate item parameters before computing DIF. Its two-sided null hypothesis is
$H_{0}\text{: }\beta-\beta_{0}=0,$ (4)
where $\beta$ denotes a population parameter such as a linear effect of an explanatory variable about the group membership in a model. The test statistic, in the form of a $z$ score that approximately follows a standard normal distribution, is
$z=\dfrac{\beta-\beta_{0}}{SE},$ (5)
where $SE$ is the standard error.
Specifically, Lord’s Wald test statistic is derived from the IRT-based item difficulty parameter from two groups for a studied item $i$, written as Langer [35]
$\hat{B}=\dfrac{\hat{b}_{F}-\hat{b}_{R}}{\sqrt{\mbox{var}(\hat{b}_{F})+\mbox{var}(\hat{b}_{R})}},$ (6)
where
$\hat{b}_{F}$ is the maximum likelihood estimate of the item difficulty parameter in the focal group;
$\hat{b}_{R}$ is the maximum likelihood estimate of the item difficulty parameter in the reference group;
$\mbox{var}(\hat{b}_{F})$ and $\mbox{var}(\hat{b}_{R})$ denote the sample variances of $\hat{b}_{F}$ and $\hat{b}_{R}$.
(6) is a special materialization of (5) for deriving the $z$ score, where the numerator is the tested parameter value and the denominator is the true standard error $SE$.
Lord extended (6) to include the IRT-based item discrimination and pseudo-guessing parameters for two groups. The hypothesis about the joint between-group difference derived from multiple IRT parameters is tested using a chi-square statistic expressed as
$\chi^{2}_{Lord}=\boldsymbol{\hat{\gamma}^{\prime}}\boldsymbol{\Sigma}^{-1}\boldsymbol{\hat{\gamma}},$
where
$\boldsymbol{\hat{\gamma}^{\prime}}$ is $\begin{bmatrix}\hat{a}_{F}-\hat{a}_{R}&\hat{b}_{F}-\hat{b}_{R}\end{bmatrix}$ when considering item discrimination together with item difficulty;
it is $\begin{bmatrix}\hat{a}_{F}-\hat{a}_{R}&\hat{c}_{F}-\hat{c}_{R}\end{bmatrix}$ when considering item discrimination together with item pseudo-guessing;
$\boldsymbol{\Sigma}^{-1}$ is the variance-covariance matrix of $\hat{\gamma}$.
Although Lord’s Wald test has not often been used for DIF detection in psychometric practice, it has been revived in modern DIF method discussions by Kim et al. [56], Langer [35], and Woods et al. [46]. Kim et al. [56] extended the method for DIF analyses with three or more groups. Langer [35] argued that the poor performance of Lord’s Wald test was partly due to its original item parameter estimation method. Woods et al. [46] compared improved versions of Lord’s Wald test.
As mentioned earlier, we first need to estimate the item parameters of this selected item when running Lord’s Wald test. It is important to highlight that estimating the parameters of a single item is not feasible because the corresponding IRT model would be under-identified. Moreover, the parameter estimate for an item fluctuates when varying the other items included in the analysis, owing to variations in the latent trait scale associated with each different item set. In practice, only one item set is needed to make conclusions about the DIF presence; this item set is chosen because it is deemed most appropriate to help describe the respondents.
Application in R. For this analysis, we used the Verbal Aggression data [57], made of the responses from 316 respondents to a questionnaire about verbal aggression. The response data were coded with 1 to mean that the respondent would want to respond to a frustrating situation aggressively, and with 0 to mean otherwise. To show the impact of the item set on Lord’s Wald test statistic, we explored this DIF analysis of our targeted item named “S1wantCurse” when using four different item sets that involved this item in addition to one, two, three other items, and all the other items in the Verbal data set, respectively. All these items were calibrated using the one-parameter item response model for parsimony. The selected items are specified below, for each of the item sets labeled with A, B, C, and D.
Table 7: Selected items in the three item sets.
Item Set N. of Items Item Names
A 2 S1wantCurse, S1WantScold.
B 3 S1wantCurse, S1WantScold,
S1WantShout.
C 4 S1wantCurse, S1WantScold,
S1WantShout, S2WantCurse.
D 24 S1wantCurse, S1WantScold,
S1WantShout, S2WantCurse,
… S4DoShout.
As we expected, our Lord’s Wald test results varied when using the three item sets, because the item parameter estimates varied when using a different item set. When analyzing the targeted item “S1wantCurse” using each of the item sets A, B and C, we found Lord’s test statistic is 0.0006, 0.0002, 0.0246, and 1.3724, respectively. All these statistics are not statistically significant at the 0.05 level. They suggest that no DIF exists between the two gender groups treated as the reference group and the focal group in expressing the tendency to respond aggressively to a frustrating situation, concerning their answers to the survey item “S1wantCurse.” We used four item sets here only to demonstrate that the composition of the item set has an impact on the DIF results.
Table 8: Lord’s Wald test results.
Item $\chi^{2}_{Lord}$ $p$-value
A 0.0006 0.9800
B 0.0002 0.9899
C 0.0246 0.8753
D 1.3724 0.2414
  • ${}^{*}p<0.05$; ${}^{**}p<0.01$; ${}^{***}p<0.001$.
The following R code demonstrates the use of an existing R function difLord() to complete Lord’s Wald test for DIF analysis of the targeted item “S1wantCurse.” The full item set (composed by 24 items) are used for this analysis, and they are calibrated using the one-parameter item response model:
1library(difR)
2data("verbal")
3
4d4 = verbal[, c(1:24, 26)]
5res4 <- difLord(Data = d4, group = "Gender", focal.name = 1, model = "1PL")
6
7res4
8#// display the test statistic
Appendix C Selected Categorical Variables from the Codebook
Table 9: The item response variable “ADDPREV.”
Code Description
1 Yes
2 No
85 BAD DATA Logically assigned
94 DON’T KNOW
97 REFUSED
98 BLANK (NO ANSWER)
99 LEGITIMATE SKIP
Table 10: The group membership variable “PDEN10.”
Code Description
1 Segment in a CBSA with 1 million or more persons
2 Segment in a CBSA with fewer than 1 million persons
3 Segment not in a CBSA
Table 11: The level-2 variable “IREDUHIGHST2.”
Code Description
1 Fifth grade or less grade completed
2 Sixth grade completed
3 Seventh grade completed
4 Eighth grade completed
5 Ninth grade completed
6 Tenth grade completed
7 Eleventh or Twelfth grade completed, no diploma
8 High school diploma/GED
9 Some college credit, but no degree
10 Associate’s degree (for example, AA, AS)
11 College graduate or higher